how to solve 12y-15=-y-37
Answer by algebrahouse.com(1079)

12y - 15 = -y - 37
13y - 15 = -37 {added y to both sides}
13y = -22 {added 15 to both sides}
y = -22/13 {divided both sides by 13}
www.algebrahouse.com